OpenStudy (anonymous):

One to one function h is defined by h(x)=(3x-4)/(-2x+9). Find h^1, the inverse of h. Then give the domain and range of h^1 using interval notation. h(x)=(3x-4)/(-2x+9) Inverse: h^-1(x)= (9x+4)/(2x+4) Don't know how to do the domain and range. Think domain is (-infinity, 9) union (4, infinity) and range is (negative infinity, infinity). Been on this all day- can't find anything or anyone that makes sense... please help

OpenStudy (anonymous):

Inverse: h^-1(x)= (9x+4)/(2x+4) Don't know how to do the domain and range. Think domain is (-infinity, 9) union (4, infinity) and range is (negative infinity, infinity). Been on this all day- can't find anything or anyone that makes sense... please help

OpenStudy (anonymous):

domain of \(h\) is all real numbers except \(\frac{9}{2}\) where the denominator is zero

OpenStudy (anonymous):

inverse is either a mistake or a typo it is \[h^{-1}(x)=\frac{9x+4}{2x+3}\]

OpenStudy (anonymous):

domain of \(h^{-1}\) is all real numbers except \(-\frac{3}{2}\) which is not surprising , because the range of \(h\) is all real numbers except \(-\frac{3}{2}\)

OpenStudy (anonymous):

It has to be in interval notation. That's my problem- it's an equations where it's one to one. I've inversed the original equation, but can't find out how to put it into interval form for the domain and range. I would name my first born satellite73 if you get this one ;)

OpenStudy (anonymous):

original was h(x) = 3x-4 ______ -2x+9

OpenStudy (anonymous):

interval notation for all real numbers except \(\frac{9}{2}\) is the silly looking \[(-\infty, \frac{9}{2})\cup (\frac{9}{2},\infty)\]

OpenStudy (anonymous):

all you need to do for the domain is to set the denominator equal to zero and solve for \(x\) then say "all numbers but that one"

OpenStudy (anonymous):

OMG you're my hero!

OpenStudy (anonymous):

somehow you made a tiny mistake on the inverse, it is not what you wrote, it is the one i wrote above

OpenStudy (anonymous):

and clearly \(\frac{3x-4}{-2x+9}\) cannot be \(-\frac{3}{2}\) if you look at it carefully you will see why

OpenStudy (anonymous):

yes i did the 4 was supposed to be a 3 in the denominator And the range

OpenStudy (anonymous):

domain of \(h\) is the range of \(h^{-1}\) and vice versa

OpenStudy (anonymous):

gotta run, good luck

OpenStudy (anonymous):

thank you!